# Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave Background

### v1

@@ Abstract
The formation and evolution of supermassive black holes (SMBHs) remains an open question in the field of modern cosmology. The detection of nanohertz (n-Hz) gravitational waves via pulsar timing arrays (PTAs) in the form of individual events and the stochastic gravitational wave background (SGWB) offers a promising avenue for studying SMBH formation across cosmic time, with SGWB signal being the immediately detectable signal with the currently accessible telescope sensitivities. By connecting the galaxy properties in the large scale (Gpc scale) cosmological simulation such as MICECAT with the small scale ($\sim$ Mpc scale) galaxy simulations from ROMULUS, we show that different scenarios of galaxy–SMBHs evolution with redshift leads to a frequency-dependent spatial anisotropy in the SGWB signal. The presence of slow evolution of the SMBHs in the Universe leads to a pronounced blue anisotropic spectrum of the SGWB. In contrast, if SMBHs grow faster in the Universe in lighter galaxies, the frequency-dependent spatial anisotropy exhibits a more flattened anisotropic spectrum. This additional aspect of the SGWB signal on top of the monopole SGWB signal, can give insight on how the SMBHs form in the high redshift Universe and its interplay with the galaxy formation from future measurements.
@@
pubyear: 2024pagerange: Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave BackgroundReferences
@@ 1 Introduction
The groundbreaking observation of gravitational waves (GW) by the LIGO-Virgo-KAGRA (LVK) collaboration (Abbott et al., 2016) has opened up new possibilities for understanding the complex nature of our Universe. The LVK collaboration has been discovering compact binary mergers, which include binary black holes (BBHs), binary neutron stars (BNSs), and neutron star-black holes (NSBHs) mergers, since 2015. (Abbott et al., 2021). Pulsar Timing Array (PTA) (Verbiest et al., 2022; Manchester, 2013; Burke-Spolaor et al., 2019) operating in the $10^{-9}-10^{-7}$ Hz range, promises to complement these findings and provide an additional avenue for studying the undiscovered domains of the cosmos. The PTA consortium, which includes the North American Observatory for Gravitational Waves (NANOGrav; McLaughlin (2013)), European PTA (EPTA; Desvignes et al. (2016)), Parkes PTA (PPTA; Manchester et al. (2013)), Indian PTA (InPTA; Joshi et al. (2018)), Chinese PTA (CPTA; Xu et al. (2023)) and MeerKAT (Bailes et al., 2020), search for the gravitational wave by accurately measuring the pulse arrival times of a set of pulsars. Each group observes a set of pulsars once every few weeks for periods of years. Recently, the PTA collaborations NANOGrav (Agazie et al., 2023a), EPTA+InPTA (Antoniadis et al., 2023), PPTA (Zic et al., 2023), and CPTA (Xu et al., 2023) have announced the evidence of a stochastic gravitational wave background (SGWB).
@@
The primary source of SGWB is believed to be the population of supermassive black hole binaries (SMBHBs) in the mass range $\sim$ $10^{7}-10^{10}M_{\odot}$ apart from cosmological sources (Ellis et al., 2024; Madge et al., 2023). These SMBHs are believed to have formed from smaller seed BHs. The seed BHs grow by the hierarchical mergers and accretion of surrounding matter (Merloni & Heinz, 2008; Cattaneo et al., 2009; Volonteri & Natarajan, 2009; Jahnke & Macciò, 2011; Ricarte et al., 2019). The history of mergers of SMBHs is highly dependent on the initial seed BHs and the environment in which these BHs evolve. Further, the environment surrounding the host galaxy influences both the growth of the SMBHs and the host galaxy through accretion and merger. This suggests coevolution of both the galaxy and the central SMBHs. There is strong evidence for the correlation between the masses of SMBHs and their host galactic properties from simulations and observational studies (Häring & Rix, 2004; Jahnke & Macciò, 2011; Beifiori et al., 2012; McConnell & Ma, 2013; Kormendy & Ho, 2013; Reines & Volonteri, 2015; Habouzit et al., 2021). Therefore, to understand the evolutionary trajectories of both central BHs and galaxies, it is essential to conduct a study that involves the simultaneous examination of SMBHBs and their host galaxies. This can help us address various questions, such as the formation, evolution, and types of galaxies favorable for hosting SMBHBs, among others. Moreover, the detection of SMBHs at the high redshift Universe by the James Webb Space Telescope (JWST) has questioned the scenarios of their formation and evolution in the Universe (Übler et al., 2023; Larson et al., 2023; Harikane et al., 2023).
@@
In this regard, the study of the GW from the SMBHBs can be very crucial. Future PTA experiments are expected to detect signals from individual low redshift SMBHBs. Apart from the individual sources at low redshifts, the SGWB from unresolved sources at higher redshifts can provide crucial insights into the population and evolution of SMBHBs across cosmic time. The shape and strength of the SGWB spectrum depend on the population of the SMBHBs and their environment. When two galaxies hosting SMBHs merge, an SMBHB is formed. These binaries lose energy to the environment through mechanisms such as dynamical friction, stellar loss-cone scattering, and viscous drag, in addition to gravitational wave (GW) radiation (Sampson et al., 2015; Kelley et al., 2017; Chen et al., 2017; Izquierdo-Villalba et al., 2022). The evolution of SMBHBs driven solely by GW emission predicts a power-law GW background spectrum. However, the spectrum of the SGWB is expected to vary significantly depending on the environment in which these binaries evolve (Saeedzadeh et al., 2023a). Environmental effects become dominant at larger separations, i.e., when the binary emits at lower frequencies, thereby influencing the shape and steepness of the spectra. Consequently, the spectrum of the SGWB can provide valuable insights into the environment in which SMBHBs are formed. In addition to the sky-averaged SGWB, the SGWB is also expected to exhibit spatial anisotropy.
In this work, we demonstrate how the frequency-dependent spatial anisotropy of the SGWB can provide valuable information about the SMBHB population and its evolution. Throughout the paper, we make the simplistic assumption of circular orbits for the SMBHBs and disregard all other potential sources of n-Hz gravitational waves, apart from SMBHBs. The presence of eccentricity in SMBHBs has been studied in the past (Armitage & Natarajan, 2005; Iwasawa et al., 2011; Bonetti et al., 2020; Gualandris et al., 2022). In this analysis, we primarily focus on the large-scale anisotropy in the n-Hz GW signal and its dependence on the cosmic evolution of SMBHs. Therefore, the impact of eccentricity is not important. The paper is organized as follows: in Sec. 2, we discuss the motivation behind the present work; in Sec. 3, we review the formation and evolution of SMBHs and SMBHBs; in Sec. 4, we discuss the cross-matching of the small-scale effect to the large scale distribution of galaxies; in Sec. 5, we briefly discuss the gravitational waves emitted by SMBHBs. Then, in Sec. 6, we describe the SGWB and the simulation used to generate the SGWB signal, along with its results. Finally, in Sec. 8, we discuss the conclusions and future prospects.
2 Motivation behind the work
The formation and evolution of SMBHs are among the leading mysteries in cosmology today (Volonteri, 2007; Jahnke & Macciò, 2011; Latif & Ferrara, 2016; Woods et al., 2019; Smith & Bromm, 2019; Sassano et al., 2021). While various models have been proposed, none have convincingly accounted for the existence of SMBHs at high redshifts. (Volonteri, 2010; Latif & Ferrara, 2016; Volonteri et al., 2021; Sassano et al., 2021). The complete understanding of the formation of SMBHs requires knowledge of the SMBH population and its evolution with redshift, along with coevolution with the host galaxies. In this context, the detection of n-Hz gravitational waves using PTA can provide a novel avenue to study the formation of SMBHs. PTAs can measure the SGWB power spectrum along with its large-scale anisotropy. The measurement of the anisotropy will provide complementary information on top of the isotropic SGWB.
@@ 2 Motivation behind the work
The motivation behind the work is to understand imprints of the evolution of SMBHs across redshift and their mergers on the large angular scale anisotropies in the SGWB as a function of its spectra, as shown in the schematic diagram Fig. 1. This figure shows that the study of the large angular anisotropies (about a degree-scale or larger; Janssen et al. (2015); Ali-Haïmoud et al. (2021); Pol et al. (2022); Sato-Polito & Kamionkowski (2023)) as a function of GW frequencies which are accessible from the current and next generation radio antennas, can make it possible to explore the formation and evolution of the SMBHBs. One of the observable quantities from the n-Hz signal is the sky-averaged SGWB signal $\Omega_{\rm GW}(f)$ and its spatial fluctuations in the spherical harmonic domain denoted by $C_{l}(f)$ at large scales (lower valuer of $l$). However, one of the major challenges in predicting the spatial fluctuations at large scales arising from the mergers of the SMBHBs across cosmic redshifts based on cosmological simulations is the involvement of large dynamical length scales from Gpc-scale to sub-parsec scale.
The SGWB signal in the n-Hz frequency range, measurable at large angular scales from sources integrated over the observable Universe, arises from the coalescence of SMBHBs at sub-parsec scales. The dynamics of SMBHBs are influenced by small-scale processes such as environmental hardening and accretion, which operate at scales ranging from sub-parsec to kiloparsec scales (Chen et al., 2017; Kelley et al., 2017; Izquierdo-Villalba et al., 2022; Saeedzadeh et al., 2023a). Resolving these binaries at such small scales, along with capturing the large-scale cosmic evolution of the Universe, is extremely challenging due to the large dynamical length scale in the problem, ranging from Gpc scale (cosmological scale) to sub-pc scale (separation between the SMBHBs emitting n-Hz). Therefore, it is crucial to incorporate the small-scale physics of SMBHB dynamics into the large cosmological scales in order to predict the impact of the redshift evolution of the SMBHBs on the n-Hz SGWB signal.
In this work, we investigate the impact of the cosmic evolution of SMBHBs on the n-Hz SGWB signal by using a large scale to small scale adaptive technique as shown by a schematic diagram in Fig. 2. In this technique, we combine physics in three length scales, Gpc scale, Mpc scale, and sub-pc scale. At the Gpc scale, we utilize the cosmological simulation MICECAT to explore the cosmic evolution of galaxies. At the Mpc scale, we employ the ROMULUS simulation to investigate the host galaxy properties of the SMBHBs. Finally, at the sub-pc scale, we use an analytical prescription to evolve the coalescence of the binaries, aiming to predict the distinguishable signatures of cosmic SMBHB evolution on the n-Hz SGWB signal. This multiscale adaptive technique enables the connection of physics across scales, facilitating an understanding of the interplay between galaxy and SMBH evolution. By using different analytical relations to connect the SMBHBs mass and merger rate evolution with host galaxy properties, we can explore and predict the observable quantities such as the sky average SGWB power spectrum $\Omega_{\rm GW}(f)$ and its spatial fluctuations power spectrum as a function of GW frequencies $C_{l}(f)$. These predictions can be tested using n-Hz GW observations for a wide range of astrophysical models.
Figure 1: A schematic diagram explaining the coevolution of the SMBH and host galaxy and its consequence on the anisotropy of the SGWB. The left panel and the right panel represent the coevolution of the SMBHs and the host galaxies at low and high redshifts, respectively.
@@ 3 Formation and Evolution of SMBHB
In our current understanding, SMBHs are believed to form through the accretion of matter and the hierarchical merger of the initial seed BHs (Volonteri et al., 2003; Merloni & Heinz, 2008; Volonteri, 2010; Woods et al., 2019; Chen et al., 2020; Wirth & Bekki, 2020; Ni et al., 2022). The precise process leading to the creation of these seeds is still a subject of ongoing research, but they represent the foundational stage for SMBHs. Some potential sources of seeds for SMBHs include remnants of the first generation of stars, massive primordial black holes, and the collapse of massive stars created in runaway collisions in dense stellar clusters (Volonteri & Natarajan, 2009; Booth & Schaye, 2009; Alexander & Natarajan, 2014; Latif & Ferrara, 2016; Pacucci et al., 2017; Ricarte & Natarajan, 2018; Sassano et al., 2021; Volonteri et al., 2021; Natarajan et al., 2023). In the standard cosmological hierarchical structure formation model, massive halos experience multiple mergers during their lifetimes (Volonteri et al., 2003; Rees & Volonteri, 2006; Ni et al., 2022). The seed BHs in early halos grow by accreting halo gas and hierarchical mergers. These BHs sink to the center of the halo due to dynamical friction and form the center of the galaxies (Volonteri, 2010; Smith & Bromm, 2019; Woods et al., 2019). These galaxies further merge to give binary BHs. Once a binary is formed, its evolution is driven by environmental effects. The three main environmental processes are dynamical friction, stellar loss cone scattering, and viscous drag. These effects are effective at different separations of the binaries. The dynamical friction, which is the retarding force experienced by the SMBH as it moves through a bulk of lighter particles, is effective at above-parsec separation (Dosopoulou & Antonini, 2017; Kelley et al., 2017). At sub-parsec separation until $\sim 10^{-2}$ pc, binary hardening is dominated by the scattering of stars by the SMBH, known as stellar loss cone scattering (Vasiliev et al., 2015; Kelley et al., 2017). Viscous hardening by circumbinary disc drives the BHs below the separation of around $10^{-2}$ before the GW emission becomes dominant at a very small separation. Although the fundamental mechanism that gives rise to the evolution of the SMBH may be known to some extent, we still do not comprehend many characteristics and phenomena associated with the SMBH evolution. Some of the most important questions in this context are listed below.
@@
    What kind of galaxies host the SMBHBs?: Several simulations, as well as observational analyses, indicate a correlation between the mass of SMBHs and the properties of their host galaxies, particularly bulge mass, stellar mass, luminosity, and velocity dispersion (Cattaneo et al., 2009; Häring & Rix, 2004; Beifiori et al., 2012; Kormendy & Ho, 2013; Reines & Volonteri, 2015; McConnell & Ma, 2013; Muhamed Kozhikkal et al., 2023). This signifies the coevolution of SMBHs and their host galaxies. However, the interplay between the growth of SMBHs and galaxies is not yet well understood. Understanding the formation and evolution of both SMBHBs and their hosts requires not just the relationship at the local Universe but also its evolution with redshift (Ding et al., 2020; Muhamed Kozhikkal et al., 2023). The key questions include: Are certain types of galaxies more conducive to the formation of SMBHBs? What is the relationship between the masses of BH binaries and the observable properties of the host galaxies? How does this relationship evolve with redshift? Do black holes occupy heavier galaxies at low redshifts or higher redshifts?
    Understanding the probability of a galaxy to host a BH of a certain mass denoted as $P(M_{\rm BH},q|{\rm galaxy})$, and its redshift evolution is crucial for addressing many questions regarding the SMBH population and its evolution. Here, $M_{\rm{1}}$ and $q$ represent the primary mass of the SMBHB and mass ratio, respectively, and ’galaxy’ denotes the properties of the host galaxies such as stellar mass, luminosity, velocity dispersion, etc. In this analysis, we use the stellar mass ($M_{*}$) of the host galaxy as an indicator to model the occupation of SMBHBs in galaxies.
@@
    What was the mass of the initial seed BHs?: Understanding the mass of the initial seed BHs is pivotal for unraveling the subsequent cosmic dynamics of SMBHs. Observational evidence suggests that there are SMBH of mass $\gtrsim$ $10^{9}\rm M_{\odot}$ at z $\gtrsim$ 6 (Barth et al., 2003; Wang et al., 2021). This provides significant constraints on the possible mass of the seed. Given that the cosmic time available at that redshift is less than a billion years, the seeds must be at least a few hundred times the mass of $\rm M_{\odot}$ (Volonteri, 2010). Some of the popular scenarios for the formation of initial seed BHs include massive Population III stars, the direct collapse of massive gas clouds, the formation of massive stars through collisions in compact nuclear star clusters, and primordial BHs. Recent work by Natarajan et al. (2023) claims that recent discovery of a quasar called UHZ1 by Chandra-JWST at redshift of around 10.1 hints towards the existence of the heavy black hole seed formed from direct collapse.
    The initial mass of SMBHs directly influences the mass distribution of SMBHs as well as the correlation between SMBH mass and host galaxy properties with redshift ($P(M_{\rm BH},q|M_{*},z$)). Understanding this correlation can help us trace back the mass of SMBHs through the evolution of the host galaxy over cosmic time.
@@
    How do SMBHs evolve over cosmic time? SMBHs are believed to have grown from initial seeds through accretion and mergers, or formed from the direct collapse of the massive gas clouds. Accretion is expected to contribute significantly to the growth at all cosmic times (Merloni & Heinz, 2008; Inayoshi et al., 2022; Lin et al., 2023). The comparison of the total mass in SMBHs today with the total integrated radiative output from quasar (Soltan, 1982) suggests that the last 2 to 3 e-folds of the growth of the SMBH are dominated by radiatively efficient accretion (Volonteri, 2010). However, a comprehensive understanding of the evolution of SMBHs requires further investigation. Studying the SMBH mass-galaxy properties relation [$P(M_{\rm BH},q|M_{*},z$)] and its evolution can provide additional insights into the growth mechanisms of SMBHs. This relation is thought to be influenced by various factors, including accretion rates, galaxy mergers, feedback processes, etc. By analyzing how this correlation changes over cosmic time, we can gain a deeper understanding of the co-evolution of SMBHs and their host galaxies.
  4. 4.
@@
    SMBHBs are the byproduct of galaxy mergers. These binaries spiral inward, emitting gravitational waves towards the end of the merger. However, the binaries evolving purely by GW emission need to be much closer than a parsec distance for them to merge on a timescale smaller than the age of the Universe. This is known as the final parsec problem (Milosavljević & Merritt, 2003; Vasiliev et al., 2015; Koo et al., 2023). For the binaries to come closer and actually emit gravitational waves in the PTA sensitive range and eventually merge, the binary must undergo hardening by the environmental processes mentioned above. The timescale of hardening by various processes depends on galactic properties such as stellar and gas density, distribution of gas in the galaxy, and velocity dispersion of stars. These effects become more significant at larger separations where gravitational wave (GW) hardening is weak (Vasiliev et al., 2015; Kelley et al., 2017; Dosopoulou & Antonini, 2017).
    The efficiency of environmental hardening directly influences the observed distribution of binary separation, $P(a|M_{*})$ as well as the occupation fraction of the binary host galaxy, where ’$a$’ is the semi-major axis of the binary. A distribution with relatively more sources at lower frequencies could indicate that binaries are efficiently being transported to smaller separations. $P(a|M_{*})$ can be further factored as the coalescence rate of the binaries at a given separation ($dN_{a}/dt_{c}$) and time spent by a binary at that separation ($dt_{c}/da$). The quantity $dt_{c}/da$ depends on the environment condition in which binary evolves. The effect of the environment on $dt_{c}/da$ can be modeled as (Sampson et al., 2015; Saeedzadeh et al., 2023a)
    $\frac{\tau_{h}}{\tau_{GW}}(f_{r})=\Big{[}1+\beta\Big{(}\frac{f_{r}}{f_{t}}\Big{)}^{-\kappa}\Big{]}^{-\gamma},$ (1)
    where $\tau_{h}(f_{r})$ and ${\tau_{GW}}(f_{r})$ represent the environmental hardening time and the GW hardening time of the binary respectively, $f_{t}$ represents the transition frequency above which the GW emission becomes dominant over the environmental effect. $\kappa$, $\gamma$, and $\beta$ are parameters that depend on the environmental properties of the host galaxy like stellar mass, gas density, stellar density, velocity dispersion, etc. Here, $\kappa$ governs the spectral dependence of the environmental effects, $\gamma$ controls the overall shape of the environmental effects, and $\beta$ governs the overall strength of the environmental effect in comparison to the hardening effect solely due to gravitational waves. We refer to Saeedzadeh et al. (2023a) for a detailed description of the parameters and their dependence on the galactic properties.

### v2

@@ Abstract
The formation and evolution of supermassive black holes (SMBHs) remains an open question in the field of modern cosmology. The detection of nanohertz (n-Hz) gravitational waves via pulsar timing arrays (PTAs) in the form of individual events and the stochastic gravitational wave background (SGWB) offers a promising avenue for studying SMBH evolution across cosmic time, with SGWB signal being the immediately detectable signal with the currently accessible telescope sensitivities. By connecting the galaxy properties in the large scale (Gpc scale) cosmological simulation such as MICECAT with the small scale ($\sim$ Mpc scale) galaxy simulations from ROMULUS, we show that different scenarios of galaxy–SMBH evolution with redshift leads to a frequency-dependent spatial anisotropy in the SGWB signal. The presence of slow evolution of the SMBHs in the Universe leads to a pronounced blue anisotropic spectrum of the SGWB. In contrast, if SMBHs grow faster in the Universe in lighter galaxies, the frequency-dependent spatial anisotropy exhibits a more flattened anisotropic spectrum. This additional aspect of the SGWB signal on top of the monopole SGWB signal, can give insight on how the SMBHs form in the high redshift Universe and its interplay with the galaxy formation from future measurements.
@@
pubyear: 2024pagerange: Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave Background11
@@ 1 Introduction
The groundbreaking observation of gravitational waves (GW) by the LIGO-Virgo-KAGRA (LVK) collaboration (Abbott et al., 2016) has opened up new possibilities for understanding the complex nature of our Universe. The LVK collaboration has been discovering compact binary mergers, which include binary black holes (BBHs), binary neutron stars (BNSs), and neutron star-black holes (NSBHs) mergers, since 2015. (Abbott et al., 2021). Pulsar Timing Array (PTA) (Verbiest et al., 2022; Manchester, 2013; Burke-Spolaor et al., 2019) operating in the $10^{-9}-10^{-7}$ Hz range promises to complement these findings and provide an additional avenue for studying the undiscovered domains of the Universe. The PTA consortia, which include the North American Observatory for Gravitational Waves (NANOGrav; McLaughlin (2013)), European PTA (EPTA; Desvignes et al. (2016)), Parkes PTA (PPTA; Manchester et al. (2013)), Indian PTA (InPTA; Joshi et al. (2018)), Chinese PTA (CPTA; Xu et al. (2023)) and MeerKAT (Bailes et al., 2020), search for the gravitational wave by accurately measuring the pulse arrival times of a set of pulsars. Each group observes a set of pulsars once every few weeks for periods of years. Recently, the PTA collaborations NANOGrav (Agazie et al., 2023a), EPTA+InPTA (Antoniadis et al., 2023), PPTA (Zic et al., 2023), and CPTA (Xu et al., 2023) have announced evidence of a stochastic gravitational wave background (SGWB).
@@
The primary source of SGWB is believed to be the population of supermassive black hole binaries (SMBHBs) in the mass range $\sim$ $10^{7}-10^{10}M_{\odot}$ apart from cosmological sources (Burke-Spolaor et al., 2019). These SMBHs are believed to have formed from smaller seed BHs. The seed BHs grew by the hierarchical mergers and accretion of surrounding matter (Volonteri, 2007; Merloni & Heinz, 2008; Volonteri & Natarajan, 2009; Jahnke & Macciò, 2011; Ricarte et al., 2019). The merger history of SMBHs is highly dependent on the environment in which these BHs evolve. The environment surrounding the host galaxy influences both the growth of the SMBH and the host galaxy. This suggests the coevolution of both the galaxies and their central SMBHs. There is strong evidence for the correlation between the masses of SMBHs and their host galaxy properties from simulations and observational studies (Häring & Rix, 2004; Jahnke & Macciò, 2011; Beifiori et al., 2012; McConnell & Ma, 2013; Kormendy & Ho, 2013; Reines & Volonteri, 2015; Habouzit et al., 2021). Therefore, to understand the evolutionary trajectories of both SMBHs and host galaxies, it is essential to conduct a study that involves the simultaneous examination of SMBHBs and their host galaxies. This can help us address various questions, such as the formation and evolution of SMBHBs, and types of galaxies favorable for hosting these binaries.
@@
In this context, the study of the GWs from the SMBHBs can be very crucial. Future PTA experiments are expected to detect signals from individual low redshift SMBHBs. Apart from the individual sources at low redshifts, the SGWB from unresolved sources at higher redshifts can provide crucial insights into the population and evolution of SMBHBs across cosmic time. The shape and strength of the SGWB spectrum depend on the population of the SMBHBs and their environment. When two galaxies hosting SMBHs merge, an SMBHB is formed. These binaries lose energy to the environment through mechanisms such as dynamical friction, stellar loss-cone scattering, and viscous drag, in addition to gravitational wave (GW) radiation (Sampson et al., 2015; Kelley et al., 2017; Chen et al., 2017; Izquierdo-Villalba et al., 2022; Bromley et al., 2023). The evolution of SMBHBs driven solely by GW emission predicts a power-law GW background spectrum. However, the spectrum of the SGWB is expected to vary significantly depending on the environment in which these binaries evolve (Saeedzadeh et al., 2023a). Environmental effects become dominant at larger separations, i.e., when the binary emits at lower frequencies, thereby influencing the shape and steepness of the spectra. Consequently, the spectrum of the SGWB can provide valuable insights into the environment in which SMBHBs are formed. In addition to the isotropic SGWB, the SGWB is also expected to exhibit spatial anisotropy (Sato-Polito & Kamionkowski, 2024; Mingarelli et al., 2013; Gardiner et al., 2024). The measurement of the anisotropy will provide complementary information on top of the isotropic SGWB.
@@
In this work, we demonstrate how we can study the imprints of the evolution of SMBHs across redshift on the large angular scale anisotropies in the SGWB as a function of its spectrum. This figure shows that the study of the large angular anisotropies (about a degree-scale or larger; Janssen et al. (2015); Ali-Haïmoud et al. (2021); Pol et al. (2022); Sato-Polito & Kamionkowski (2024)) as a function of GW frequencies which are accessible from the current and next generation radio antennas, can make it possible to explore the formation and evolution of the SMBHBs. One of the observable quantities from the n-Hz signal is the sky-averaged SGWB signal $\Omega_{\rm gw}(f)$ and its spatial fluctuations in the spherical harmonic space denoted by $C_{\ell}(f)$ at large scales (lower values of $\ell$). However, one of the major challenges in predicting the spatial fluctuations of the signal at large scales arising from the mergers of the SMBHBs across cosmic redshifts based on cosmological simulations is the involvement of large dynamical length scales from the Gpc scale to sub-parsec scale. Therefore, in this work, we investigate the impact of the cosmic evolution of SMBHBs on the n-Hz SGWB signal by using a large scale to small scale adaptive technique as shown by a schematic diagram in Fig. 1. In this technique, we combine the results of simulations with analytical prescriptions across three length scales: Gpc scale, Mpc scale, and sub-pc scale. The detailed methodology of this technique will be discussed in Sec. 3.
Throughout the paper, we make the simplistic assumption of circular orbits for the SMBHBs and disregard all other potential sources of n-Hz gravitational waves, apart from SMBHBs. The presence of eccentricity in SMBHBs has been studied in the past (Armitage & Natarajan, 2005; Iwasawa et al., 2011; Bonetti et al., 2020; Gualandris et al., 2022). In this analysis, we primarily focus on the large-scale anisotropy in the n-Hz GW signal and its dependence on the cosmic evolution of SMBHs. Therefore, the impact of eccentricity is not important. The paper is organized as follows: in Sec. 2, we review the formation and evolution of SMBHs and SMBHBs; in Sec. 3, we discuss the cross-matching of the small-scale effect to the large scale distribution of the galaxies; in Sec. 4, we briefly discuss the SGWB from SMBHB population; in Sec. 5, we describe the simulation used to generate the SGWB signal. Then, in Sec. 6 we discuss the results of the simulation. Finally, in Sec. 7, we discuss the conclusions and future prospects.
@@ 2 Formation and Evolution of SMBHB
In our current understanding, SMBHs are believed to form through the accretion of matter and the hierarchical merger of the initial seed BHs (Volonteri et al., 2003; Merloni & Heinz, 2008; Volonteri, 2010; Woods et al., 2019; Chen et al., 2020; Wirth & Bekki, 2020; Ni et al., 2022). The precise process leading to the formation of these seeds is still a subject of ongoing research, but they represent the foundational stage for SMBHs. The seed BHs in early halos grow by accreting halo gas and by hierarchical mergers. These BHs sink to the center of the halo due to dynamical friction and form the center of the galaxies (Volonteri, 2010; Smith & Bromm, 2019; Woods et al., 2019). These galaxies further merge to give binary BHs. Although the fundamental mechanism that gives rise to the evolution of the SMBH may be known to some extent, we still do not comprehend many aspects and phenomena associated with the SMBH evolution. Some of the most important questions in this context are listed below.
@@
    What kind of galaxies host the SMBHBs?:
    Several simulations, as well as observational analyses, indicate a correlation between the mass of SMBHs and the properties of their host galaxies, particularly bulge mass, stellar mass, luminosity, and velocity dispersion (Cattaneo et al., 2009; Häring & Rix, 2004; Beifiori et al., 2012; Kormendy & Ho, 2013; Reines & Volonteri, 2015; McConnell & Ma, 2013; Muhamed Kozhikkal et al., 2023). This signifies the coevolution of SMBHs and their host galaxies. However, the interplay between the growth of SMBHs and galaxies is not yet well understood. Understanding the formation and evolution of both SMBHBs and their hosts requires not just their relationship at the local Universe but also its evolution with redshift (Ding et al., 2020; Muhamed Kozhikkal et al., 2023).
    The probability of a galaxy to host a SMBHB of a certain mass denoted as $P(M_{\rm BH},q|{\rm\texttt{galaxy}})$, and its redshift evolution is crucial for addressing many questions regarding the SMBH population and its evolution. Here, $M_{\rm{1}}$ and $q$ represent the primary mass of the SMBHB and mass ratio, respectively, and ’galaxy’ denotes the properties of the host galaxies such as stellar mass, luminosity, velocity dispersion, etc. In this analysis, we use the stellar mass ($M_{*}$) of the host galaxy as an indicator to model the occupation of SMBHBs in galaxies.
@@
    How did SMBHs form and evolve over cosmic time?
    The potential sources of the seed BHs include remnants of the first generation of stars, massive primordial black holes, and the collapse of massive stars formed in runaway collisions within dense stellar clusters (Latif & Ferrara, 2016; Sassano et al., 2021; Volonteri et al., 2021). These seed BHs co-evolve with their host galaxies accreting matter from their surrounding. Moreover, galaxy mergers, which were more common in the earlier, denser universe, contributed to the growth of SMBHs. Over billions of years, these processes allowed SMBHs to reach masses ranging from millions to billions of solar masses.
    Studying the SMBH mass-galaxy stellar mass relation [$P(M_{\rm BH},q|M_{*},z$)] and its evolution can potentially provide insights into the intricacies of the growth mechanisms of SMBHs. This relation is thought to be influenced by various factors, including accretion rates, galaxy mergers, feedback processes, etc. By analyzing how this correlation changes over cosmic time, we can gain a deeper understanding of the formation and evolution of SMBHs.
@@
    SMBHBs are the byproduct of galaxy mergers. These binaries spiral inward under the influence of their surroundings, such as dynamical friction and viscous drag. These processes are known to be effective until binary separation is greater than a parsec. However, the GW radiation becomes an effective source of orbital energy decay when the binary separation is much below parsec (around $10^{-3}-10^{-2}$ parsec). The challenge of how these binaries can come close enough to merge within the Hubble time is known as the final parsec problem (Milosavljević & Merritt, 2003; Vasiliev et al., 2015; Koo et al., 2023). For the binaries to emit GWs in the PTA-sensitive range and eventually merge, the environmental processes mentioned above must be effective enough to bring the binaries to the separation where the GW radiation becomes an efficient process of energy loss.
    The efficiency of environmental hardening directly influences the observed distribution of binary separation, $P(a|M_{*})$ as well as the occupation fraction of the binary host galaxy, where ’$a$’ is the semi-major axis of the binary. A distribution with relatively more sources emitting at higher frequencies could indicate that binaries are efficiently being transported to smaller separations. $P(a|M_{*})$ can be further factored as the coalescence rate of the binaries at a given separation ($dN_{a}/dt_{c}$) and time spent by a binary at that separation ($dt_{c}/da$). The quantity $dt_{c}/da$ depends on the environment condition in which binary evolves. The effect of the environment on $dt_{c}/da$ can be modeled as (Sampson et al., 2015; Saeedzadeh et al., 2023a)
    $\frac{\tau_{\rm h}}{\tau_{\rm GW}}(f_{r})=\Big{[}1+\beta\Big{(}\frac{f_{r}}{f_{t}}\Big{)}^{-\kappa}\Big{]}^{-\gamma},$ (1)
    where $\tau_{h}(f_{\rm r})$ and ${\tau_{\rm GW}}(f_{r})$ represent the binary hardening time and the GW hardening time of the binary respectively, $f_{t}$ represents the transition frequency above which the GW emission becomes dominant over the environmental effect. $\kappa$, $\gamma$, and $\beta$ are parameters that depend on the environmental properties of the host galaxy like stellar mass, gas density, stellar density, velocity dispersion, etc. Here, $\kappa$ governs the spectral dependence of the environmental effects, $\gamma$ controls the overall shape of the environmental effects, and $\beta$ governs the overall strength of the environmental effect in comparison to the hardening effect solely due to gravitational waves. We refer to Saeedzadeh et al. (2023a) for a detailed description of the parameters and their dependence on the galactic properties.